# Generalized Entropy Calibration for Inference with Partially Observed Data: A Unified Framework

### v1

@@ 8.1 Causal inference: the LaLonde job-training study
We assess the proposed estimators by estimating the treatment effect of the National Supported Work (NSW) Demonstration, a randomized job-training program conducted in the mid-1970s (LaLonde, 1986; Dehejia and Wahba, 1999). Following the standard benchmark design, we pool the NSW experimental sample with two large observational control groups—the Panel Study of Income Dynamics (PSID) and the Current Population Survey (CPS)—into a single sample. Units are classified by a group indicator $G_{i}\in\{1,2,3,4\}$: NSW treated ($N_{1}=185$), NSW controls ($N_{2}=260$), PSID ($N_{3}=2{,}490$), and CPS ($N_{4}=15{,}992$).
The target estimand is the average treatment effect for the NSW experimental population,
$\mathrm{ATE}_{\mathrm{NSW}}\equiv\mathbb{E}\{Y(1)-Y(0)\mid\text{NSW participants}\}.$
@@
We estimate generalized propensity scores $\pi_{ig}=\Pr(G_{i}=g\mid x_{i})$ via multinomial logistic regression and define tilting weights $r_{ig}=\pi_{i,\mathrm{NSW}}/\pi_{ig}$, where $\pi_{i,\mathrm{NSW}}=\pi_{i1}+\pi_{i2}$. The covariates used for propensity score estimation are age, years of education, indicators for Black and Hispanic ethnicity, marital status, an indicator for no high school degree, and real earnings in 1974 and 1975 (re74 and re75). These are the standard covariates used in the LaLonde benchmark literature (Dehejia and Wahba, 1999). Cross-fitted outcome predictions $\hat{y}_{i}(g)$ are obtained using GAM with $K=4$ folds. For each candidate control group $g\in\{2,3,4\}$, calibration weights are constructed by solving (5.2)–(5.4) with the two-loop procedure, and the ATE is estimated as $\widehat{\mathrm{ATE}}_{g}=\bar{Y}_{1}-\hat{\theta}_{g}$, where $\bar{Y}_{1}$ is the mean outcome among NSW treated units.
Results.
@@
Table 3 reports point estimates, standard errors, and evaluation bias (EB), defined as the difference from the NSW experimental benchmark $\hat{\theta}_{\mathrm{NSW}}=1794$. Using the NSW experimental data, all estimators yield estimates close to the benchmark, as expected under randomization. Using PSID or CPS as external controls, the unweighted difference in means produces large EB (17,000-17{,}000 and 10,292-10{,}292, respectively), reflecting substantial covariate imbalance between the NSW treated group and the observational control samples. Among all weighting methods, the proposed ET estimator most closely reproduces the experimental benchmark in both external-control settings, achieving the smallest absolute EB (139-139 for PSID; 179-179 for CPS) with standard errors comparable to competing methods. For the PSID comparison, the next-best method is IPW with EB =320=-320, more than twice the absolute bias of ET. For the CPS comparison, AIPW (GAM) achieves EB =111=-111, which is slightly smaller in absolute value than ET’s 179-179, but ET has a smaller standard error ($635$ vs. $656$), so the two are broadly comparable. Overall, ET provides the most consistently accurate reproduction of the experimental benchmark across both external-control settings.
@@
Figure 5 compares weighted covariate distributions (age, education, re74, and re75) for PSID and CPS controls against the empirical distribution of the combined NSW sample. ET achieves the closest overall covariate balance across all four variables in both panels. For the PSID controls, IPW shows noticeable residual imbalance in age and re74, with the weighted CDF deviating visibly from the NSW target, while CBPS and EBCW improve upon IPW but still exhibit some discrepancy in the tails of the earnings distributions. For the CPS controls, the covariate distributions are initially closer to the NSW sample, so all weighting methods perform reasonably well, though ET and EBCW achieve the tightest alignment. These patterns are consistent with the numerical results in Table 3: methods that achieve better covariate balance tend to produce smaller evaluation bias.
@@
Table 3: Comparison of Average Treatment Effect Estimators on the NSW Experimental and PSID and CPS observational data. Notes: SE= standard error, EB = Evaluation Bias.
@@ 8.2 Semi-supervised learning: NHANES fasting glucose
We use the 2017–2018 National Health and Nutrition Examination Survey (NHANES) to estimate the association between fasting plasma glucose and cardiometabolic risk factors. NHANES is a nationally representative health survey conducted by the U.S. Centers for Disease Control and Prevention that collects detailed demographic, anthropometric, and clinical information on participants. Fasting glucose is measured only for participants attending the morning fasting examination, creating a natural semi-supervised setting: covariates are available for the full sample of N=6,230N=6{,}230 individuals, while glucose is observed for n=2,533n=2{,}533 (the labeled sample).
The covariates included in the regression model are age (years), sex (female indicator), race/ethnicity (indicators for Hispanic, non-Hispanic White, non-Hispanic Black, Asian, and Other), body mass index (BMI, kg/m2), systolic blood pressure (SBP, mmHg), and diastolic blood pressure (DBP, mmHg). These variables are standard cardiometabolic risk factors and are fully observed for all participants.
To better reflect practical semi-supervised scenarios with limited labeled data, we randomly subsample 50% of the labeled set, yielding n=1,266n=1{,}266 labeled and Nn=3,697N-n=3{,}697 unlabeled observations ($\approx$74% unlabeled). We compare the proposed ET estimator with the supervised (OLS) estimator, the density-ratio estimator (DRESS; Kawakita and Kanamori, 2013), and the projection-based estimator (PSSE; Song et al., 2024), using $K=3$ folds for cross-fitting.
Results.
Table 4 reports point estimates, standard errors, confidence interval widths (CIW), and asymptotic relative efficiency (ARE) relative to the supervised estimator. ARE values greater than one indicate that the semi-supervised estimator is more efficient (has smaller asymptotic variance) than the supervised baseline.
@@ Results.
The proposed ET estimator achieves the largest efficiency gains across methods. For the age coefficient, ET yields a standard error of $0.05$ compared with $0.06$ for the supervised estimator (ARE =1.33=1.33), while DRESS and PSSE both have SE =0.07=0.07 (ARE $\approx 0.70$), indicating that these competing methods are actually less efficient than the supervised estimator for this coefficient. A similar pattern holds for the sex coefficient: ET achieves SE =1.59=1.59 versus the supervised SE =1.99=1.99 (ARE =1.56=1.56), whereas DRESS and PSSE have SE $\approx 1.98$ with ARE barely above $1.0$. The largest efficiency gain for ET appears in the “Other race” coefficient, where ARE =2.65=2.65 corresponds to a confidence interval width of $11.59$ compared with $18.88$ for the supervised estimator—a reduction of nearly 39%.
DRESS and PSSE produce ARE values below one for several coefficients (e.g., intercept, Hispanic, SBP), meaning they are less efficient than the supervised estimator. This occurs because these methods rely on the outcome prediction model, and when the prediction is imperfect, the additional variability from incorporating unlabeled data can outweigh the efficiency gain. In contrast, ET’s debiasing constraint provides a safeguard against this phenomenon: by incorporating the propensity-score information, it ensures that the unlabeled data contribute to efficiency even when the outcome model is imperfect, consistent with the doubly robust property established in Section 4.
@@ 9 Conclusion
We have proposed a unified calibration-weighting framework for estimating parameters defined by general estimating equations with partially observed data. The method constructs weights by minimizing a generalized entropy subject to balancing constraints that incorporate cross-fitted machine-learning predictions and debiasing constraints based on a working propensity-score model. The resulting estimator is doubly robust and attains the semiparametric efficiency bound when both the outcome-regression and propensity-score models are correctly specified. A distinctive theoretical property, established in Corollary 3, is that the proposed estimator has smaller asymptotic variance than the classical AIPW estimator whenever the OR model is misspecified, because the debiasing constraint enriches the calibration space and reduces the residual variance of the estimating function after projection.
This theoretical advantage is confirmed empirically. In the causal inference simulations, HD and ET exhibit noticeably smaller variability than all competitors under outcome-model misspecification (OR2PS1), while matching AIPW under correct specification. In the missing-covariates setting, the proposed HD estimator reduces RMSE by approximately 32–36% relative to AIPW when the OR model is misspecified and the PS model is correct. In the NHANES semi-supervised application, the ET estimator achieves asymptotic relative efficiency gains of up to 2.65 over the supervised estimator, whereas competing semi-supervised methods sometimes perform worse than the supervised baseline. These results collectively demonstrate that generalized entropy calibration provides a practical and reliable approach to inference with partially observed data, particularly in settings where the outcome model may be misspecified.
Several other extensions merit investigation. First, the current framework assumes a missing-at-random mechanism; extending the approach to handle non-ignorable missingness, where the probability of observation depends on the unobserved values themselves, would substantially broaden its applicability. Second, when the dimension of the calibration function $b(O)$ is large relative to the sample size, the entropy calibration problem may become ill-conditioned; incorporating regularization into the dual objective or employing dimension-reduction techniques for the calibration constraints could address this challenge. More broadly, the proposed framework provides a general tool for inference with partially observed data and can be adapted to problems involving data integration across multiple sources.
Appendix
In this supplementary material, we provide all regularity conditions (Section A), proofs of the theoretical results stated in the main paper (Section B), and asymptotic theory for cross-fitted GEC Estimators (Section C).
Appendix A Assumptions
We organize the assumptions into three groups: conditions for the consistency of the nuisance parameters $\hat{\varphi}$ and $\hat{\lambda}$ (Assumptions 12), conditions for the linearization of the weighted estimating equation (Assumptions 34), and conditions for the joint asymptotic normality of $\hat{\theta}_{\omega}$ (Assumptions 57).
Consistency of $\hat{\varphi}$ and $\hat{\lambda}$.
Assumption 1 (PS model regularity).
The limiting function $\mathbb{E}\{\nabla\ell(\varphi)\}$ is continuously differentiable with respect to $\varphi$ on a compact set $\mathcal{G}_{\varphi}$ containing $\varphi^{*}$.
Assumption 2 (Calibration regularity).
  1. (i)
    $b_{i}$ and $O_{i}$ have compact support.
  2. (ii)
    $\mathbb{E}(\delta\,s_{i}^{*\top}s_{i}^{*})$ is nonsingular, where $s_{i}^{*}=s_{i}(\varphi^{*})$.
  3. (iii)
    The convex conjugate $F$ of $G$ is strictly convex.
Linearization.
Assumption 3 (Smoothness of limiting objectives).
The limiting functions $\mathbb{E}\{\nabla\rho_{G}(\lambda)\}$ and $\mathbb{E}\{\nabla\ell(\varphi)\}$ are differentiable, and their Hessians $\mathbb{E}\{\nabla^{2}\rho_{G}(\lambda)\}$ and $\mathbb{E}\{\nabla^{2}\ell(\varphi)\}$ are continuous on compact sets $\mathcal{G}_{\lambda}\ni\lambda^{*}$ and $\mathcal{G}_{\varphi}\ni\varphi^{*}$, respectively.
Assumption 4 (Non-degeneracy).
The matrices $\mathbb{E}\{\nabla^{2}\ell(\varphi^{*})\}$ and $\mathbb{E}\{\nabla^{2}\rho_{G}(\lambda^{*})\}$ are non-singular.
Joint asymptotic normality.
Assumption 5 (Existence and identifiability).
Let $\alpha=(\theta^{\top},\lambda^{\top},\varphi^{\top})^{\top}\in\mathbb{R}^{2q+r+1}$ and
$\hat{Q}(\alpha)\;=\;\bigl(\nabla\ell(\varphi)^{\top},\;\nabla\rho_{G}(\lambda)^{\top},\;\hat{U}^{\omega}(\theta)^{\top}\bigr)^{\top}.$
The equation $\hat{Q}(\alpha)=0$ has a unique solution $\hat{\alpha}=(\hat{\theta}^{\top},\hat{\lambda}^{\top},\hat{\varphi}^{\top})^{\top}$, and there exists a function $\mathbb{E}\{\hat{Q}(\alpha)\}$ such that $\hat{Q}(\alpha)\to\mathbb{E}\{\hat{Q}(\alpha)\}$ uniformly as $N\to\infty$ and $\mathbb{E}\{\hat{Q}(\alpha)\}=0$ has a unique solution $\alpha^{*}=(\theta^{*\top},\lambda^{*\top},\varphi^{*\top})^{\top}$.
Assumption 6 (Smoothness of the joint system).
The limiting function $\mathbb{E}\{\hat{Q}(\alpha)\}$ is differentiable and $\mathbb{E}\{\partial\hat{Q}(\alpha)/\partial\alpha\}$ is continuous on a compact set $\mathcal{G}=\mathcal{G}_{\theta}\times\mathcal{G}_{\lambda}\times\mathcal{G}_{\varphi}$ containing $\alpha^{*}$.
Assumption 7 (Non-singularity of the joint Jacobian).
The matrix $\mathbb{E}\{\partial\hat{Q}(\alpha)/\partial\alpha\}\big|_{\alpha=\alpha^{*}}$ is non-singular.
Appendix B Proofs
B.1 Proof of Lemma S.1 (Consistency of $\hat{\varphi}$ and $\hat{\lambda}$)
Lemma B.1 (Consistency).
Under Assumptions 12, $\hat{\varphi}\xrightarrow{p}\varphi^{*}$ and $\hat{\lambda}\xrightarrow{p}\lambda^{*}$, where $\varphi^{*}=\arg\max_{\varphi}E\{\ell(\varphi)\}$ and $\lambda^{*}=\arg\min_{\lambda}E\{\hat{\rho}_{G}(\lambda)\}$.
Proof.
Step 1: Consistency of $\hat{\varphi}$.  The pseudo-true value $\varphi^{*}$ minimizes the Kullback–Leibler divergence
$\mathrm{KL}(\varphi)\;=\;\mathbb{E}\!\left[\log\frac{g(\delta\mid O)}{f(\delta\mid O;\,\varphi)}\right],$
where $g(\delta\mid O)$ is the true conditional density of $\delta$ given $O$ and $f(\delta\mid O;\,\varphi)=\pi(O;\varphi)^{\delta}\{1-\pi(O;\varphi)\}^{1-\delta}$ is the working model. By Assumption 1 and Theorem 2.2 of White (1982), $\hat{\varphi}\xrightarrow{p}\varphi^{*}$.
Step 2: Uniform convergence of $\nabla\hat{\rho}_{G}$.  By the consistency of $\hat{\varphi}$ and the delta method,
$\nabla\hat{\rho}_{G}(\lambda)\;=\;\nabla\rho_{G}(\lambda)+o_{p}(1)$ (B.1)
[…]
$\sup_{\lambda\in\mathcal{G}_{\lambda}}\|\nabla\hat{\rho}_{G}(\lambda)-\nabla\rho_{G}(\lambda)\|\;=\;o_{p}(1).$ (B.2)
Step 3: Uniqueness of $\lambda^{*}$ and consistency of $\hat{\lambda}$.  Suppose $\nabla\rho_{G}(\lambda_{1})=\nabla\rho_{G}(\lambda_{2})=0$. Then
$\mathbb{E}\!\bigl[\delta\,\{g^{-1}(\lambda_{1}^{\top}s^{*})-g^{-1}(\lambda_{2}^{\top}s^{*})\}(\lambda_{1}^{\top}s^{*}-\lambda_{2}^{\top}s^{*})\bigr]=0.$
Since $g^{-1}$ is strictly increasing, the integrand is non-negative, so $\lambda_{1}^{\top}s^{*}=\lambda_{2}^{\top}s^{*}$ a.s. for $\delta=1$. Non-singularity of $\mathbb{E}(\delta\,s^{*\top}s^{*})$ (Assumption 2(ii)) implies $\lambda_{1}=\lambda_{2}$. Combined with (B.2), Theorem 5.9 of Van der Vaart (2000) gives $\hat{\lambda}\xrightarrow{p}\lambda^{*}$. ∎
B.2 Proof of Theorem 4.1 (Linearization)
Proof.
[…]
Step 1: $\sqrt{N}$-rate for $\hat{\varphi}$ and $\hat{\lambda}$.  By a mean-value expansion of $\nabla\ell(\hat{\varphi})=0$ around $\varphi^{*}$ and Assumptions 34,
$\hat{\varphi}-\varphi^{*}\;=\;-\bigl[\mathbb{E}\{\nabla^{2}\ell(\varphi^{*})\}\bigr]^{-1}\,\nabla\ell(\varphi^{*})\;+\;o_{p}(N^{-1/2}),$ (B.3)
where $\nabla\ell(\varphi)=N^{-1}\sum_{i=1}^{N}(\delta_{i}/\pi_{i}-1)\,h_{i}(\varphi)$. An analogous expansion of $\nabla\hat{\rho}_{G}(\hat{\lambda})=0$ yields
$\hat{\lambda}-\lambda^{*}\;=\;-\bigl[\mathbb{E}\{\nabla^{2}\rho_{G}(\lambda^{*})\}\bigr]^{-1}\,\nabla\rho_{G}(\lambda^{*})\;+\;o_{p}(N^{-1/2}).$ (B.4)
Step 2: Expansion of $\hat{U}^{\omega}(\theta)$ around $\lambda^{*}$.  By the mean value theorem and (B.4),
$\hat{U}^{\omega}(\theta)$ $=\frac{1}{N}\sum_{i=1}^{N}\delta_{i}\,\omega_{i}^{*}(\lambda^{*})\,U(\theta;z_{i})\;+\;\left(\frac{1}{N}\sum_{i=1}^{N}\delta_{i}\,U(\theta;z_{i})\,\frac{\partial}{\partial\lambda}\omega_{i}^{*}(\tilde{\lambda})^{\top}\right)(\hat{\lambda}-\lambda^{*})$
$=\frac{1}{N}\sum_{i=1}^{N}\delta_{i}\,\omega_{i}^{*}(\lambda^{*})\,U(\theta;z_{i})\;-\;\gamma^{*}\,\nabla\rho_{G}(\lambda^{*})\;+\;o_{p}(N^{-1/2}),$ (B.5)
where the second equality uses Randles (1982) and
$\gamma^{*}\;=\;\mathbb{E}\!\bigl[\delta\,f^{\prime}(\lambda^{*\top}S)\,U(\theta;Z)\,S^{\top}\bigr]\;\bigl\{\mathbb{E}\!\bigl[\delta\,f^{\prime}(\lambda^{*\top}S)\,S\,S^{\top}\bigr]\bigr\}^{-1},$ (B.6)
with $S=S(O;\varphi^{*})$ and $f^{\prime}=(g^{-1})^{\prime}$.
Step 3: Assembling the linearization.  Substituting the definition of $\nabla\rho_{G}(\lambda^{*})$ into (B.5) and rearranging gives
$\hat{U}^{\omega}(\theta)\;=\;\frac{1}{N}\sum_{i=1}^{N}\Bigl[\gamma^{*}\,s_{i}(\hat{\varphi})\;+\;\delta_{i}\,\omega_{i}^{*}(\lambda^{*},\hat{\varphi})\bigl\{U(\theta;z_{i})-\gamma^{*}\,s_{i}(\hat{\varphi})\bigr\}\Bigr]+o_{p}(N^{-1/2}),$
which is the claimed form $\tilde{U}^{\omega}(\lambda^{*},\hat{\varphi})+o_{p}(N^{-1/2})$. No assumption on the correctness of either the PS or OR model was used. ∎
B.3 Proof of Lemma 4.2
Proof.
Set $\lambda^{*}=(0^{\top},1)^{\top}$. Under the correct PS model, $P(\delta=1\mid O)=\pi(O;\varphi_{0})$, so
$\mathbb{E}\{\nabla\rho_{G}(\lambda^{*})\}$ $=\mathbb{E}\!\left[\delta\,g^{-1}\!\bigl(g(\pi^{-1}(O;\varphi_{0}))\bigr)\,s^{*}\;-\;s^{*}\right]=\mathbb{E}\!\left[\frac{\delta}{\pi(O;\varphi_{0})}\,s^{*}-s^{*}\right]=0,$
where the last equality follows from $\mathbb{E}(\delta\mid O)=\pi(O;\varphi_{0})$. Since the solution is unique (Lemma B.1), $\lambda^{*}=(0^{\top},1)^{\top}$ is the probability limit, giving $\omega_{i}^{*}(\lambda^{*},\varphi_{0})=\pi^{-1}(O_{i};\varphi_{0})$. ∎
B.4 Proof of Corollary 1
Proof.

### v2

@@ 8 Real-Data Applications
We conclude with two real-data applications illustrating the practical scope of the proposed framework. The goal of this section is not merely to compare numerical performance across methods, but to show how the same calibration logic applies in substantively different partially observed-data problems. The first application is a causal-inference benchmark, where one may assess performance against an experimental reference and examine covariate balance directly. The second is a semi-supervised learning problem from public health, where outcomes are observed only for a labeled subset and efficiency gains from unlabeled covariates are of primary interest.
Taken together, these examples illustrate two complementary aspects of the framework. In the causal setting, calibration helps reconstruct the target treatment comparison from observational controls while maintaining close balance on relevant covariates. In the semi-supervised setting, calibration uses predictions for the unlabeled outcomes to improve efficiency while preserving a debiasing correction tied to the observation mechanism.
@@ 8.1 Causal inference: the LaLonde job-training study
We first consider the National Supported Work (NSW) Demonstration, a canonical benchmark for causal inference with observational controls (LaLonde, 1986; Dehejia and Wahba, 1999). This application is useful for two reasons. First, the experimental NSW estimate provides a reference against which observational estimators can be compared through evaluation bias. Second, the weighting structure can be examined directly through covariate balance, making the role of calibration transparent.
@@
Following the standard benchmark design, we pool the NSW experimental sample with two large observational control groups, the Panel Study of Income Dynamics (PSID) and the Current Population Survey (CPS), into a single sample. Units are classified by a group indicator $G_{i}\in\{1,2,3,4\}$: NSW treated ($N_{1}=185$), NSW controls ($N_{2}=260$), PSID ($N_{3}=2{,}490$), and CPS ($N_{4}=15{,}992$). The target estimand is the average treatment effect for the NSW experimental population,
$ATE_{\text{NSW}}\equiv E\{Y(1)-Y(0)\mid\text{NSW participants}\}.$
@@
We estimate generalized propensity scores $\pi_{ig}=P(G_{i}=g\mid x_{i})$ via multinomial logistic regression and define tilting weights $r_{ig}=\pi_{i,\text{NSW}}/\pi_{ig}$, where $\pi_{i,\text{NSW}}=\pi_{i1}+\pi_{i2}$. The covariates used for propensity-score estimation are age, years of education, indicators for Black and Hispanic ethnicity, marital status, an indicator for no high school degree, and real earnings in 1974 and 1975 (re74 and re75), following the standard LaLonde benchmark literature. Cross-fitted outcome predictions $\hat{y}_{i}(g)$ are obtained using GAM with $K=4$ folds. For each candidate control group $g\in\{2,3,4\}$, calibration weights are constructed by solving (18)–(20) with the two-loop procedure, and the ATE is estimated as $\widehat{ATE}^{[g]}=\bar{Y}_{1}-\hat{\theta}_{g}$, where $\bar{Y}_{1}$ is the mean outcome among NSW treated units.
@@
Results. Table 4 shows that generalized entropy calibration performs competitively with established weighting and doubly robust methods in both external-control settings, while remaining close to the experimental benchmark $\hat{\theta}_{\text{NSW}}=1794$. Using the NSW experimental data alone, all estimators are close to the benchmark, as expected under randomization. Using PSID or CPS as external controls, the unweighted difference in means exhibits very large evaluation bias, reflecting substantial covariate imbalance between the NSW treated group and the observational control samples. Among the weighting estimators, ET yields one of the smallest evaluation biases in both settings: for PSID, its absolute bias is smaller than that of IPW, CBPS, EBCW, and the two AIPW implementations; for CPS, its estimate is again close to the benchmark and broadly comparable to the strongest competing approaches, with a standard error similar to or smaller than several alternatives. Overall, the pattern suggests that generalized entropy calibration provides a favorable balance between bias reduction and precision across both external-control samples.
@@
Figure 5 compares weighted covariate distributions for age, education, re74, and re75 in the PSID and CPS control samples against the empirical distribution of the combined NSW sample. The figure shows that generalized entropy calibration achieves close alignment across these key variables in both panels. For the PSID controls, the contrast is especially visible: IPW retains noticeable residual imbalance in age and re74, while CBPS and EBCW improve the fit but still leave discrepancies in parts of the earnings distributions. For the CPS controls, all methods perform more similarly because the untreated sample is already closer to the NSW sample, though ET and EBCW continue to show the closest overall agreement. These balance patterns are consistent with the evaluation-bias comparisons in Table 4: methods that achieve better covariate balance tend to exhibit smaller evaluation bias. This application therefore illustrates the causal-inference face of the proposed framework: calibration improves agreement with the experimental benchmark not simply by reweighting, but by enforcing a balance structure that better aligns the observational controls with the treated sample.
Table 4: Comparison of Average Treatment Effect Estimators on the NSW Experimental and PSID and CPS observational data.
Notes: SE= standard error, EB = Evaluation Bias.
@@ 8.2 Semi-supervised learning: NHANES fasting glucose
We next consider a semi-supervised regression problem using the 2017–2018 National Health and Nutrition Examination Survey (NHANES). Fasting plasma glucose is measured only for participants attending the morning fasting examination, whereas demographic and cardiometabolic covariates are available for the full sample. This creates a natural labeled–unlabeled structure and provides a concrete setting in which to examine whether calibration can convert information in the unlabeled covariate distribution into more efficient regression inference.
We use NHANES to estimate the association between fasting plasma glucose and cardiometabolic risk factors. The covariates in the regression model are age, sex, race/ethnicity indicators, body mass index (BMI), systolic blood pressure (SBP), and diastolic blood pressure (DBP). These variables are fully observed for all participants. Fasting glucose is available for n=2,533n=2{,}533 of the N=6,230N=6{,}230 individuals, forming the labeled sample. To mimic a practically relevant semi-supervised scenario with limited labeled data, we randomly retain only 50% of the originally labeled observations, yielding n=1,266n=1{,}266 labeled and Nn=3,697N-n=3{,}697 unlabeled units. We compare the proposed ET estimator with the supervised (OLS) estimator, the density-ratio estimator (DRESS; Kawakita and Kanamori, 2013), and the projection-based estimator (PSSE; Song et al., 2024), using $K=3$ folds for cross-fitting.
Results. Table 5 shows that generalized entropy calibration yields clear efficiency gains over the supervised estimator across several regression coefficients, often outperforming the competing semi-supervised alternatives as well. The improvement is most visible in the reduction of standard errors and confidence-interval widths, indicating that the unlabeled sample is being used effectively. In several coefficients, ET attains substantial asymptotic relative efficiency gains compared with the supervised estimator, while maintaining stable point estimates. More broadly, the NHANES analysis shows that the calibration framework is not limited to causal weighting problems: the same balance-and-debias logic can be used to strengthen regression inference when responses are available only for a subset of the sample. In this application, the practical value of the framework lies less in reproducing a benchmark than in improving efficiency while preserving a principled correction for selective outcome availability.
These two analyses highlight complementary aspects of generalized entropy calibration. In the LaLonde benchmark, the method operates as a balancing-weight procedure for causal inference and improves agreement with an experimental reference through closer covariate alignment. In NHANES, it operates as a semi-supervised estimator that converts unlabeled covariate information into more efficient regression inference. The common thread is that both applications are instances of the same partially observed-data template: a full-data estimating equation is reconstructed through calibration constraints that combine outcome-related information with a debiasing correction.
@@ 9 Conclusion
This paper develops a unified calibration-based framework for inference with partially observed data. The central idea is that augmentation can be represented implicitly through calibration: rather than adding an explicit augmentation term to an inverse-probability weighted estimating equation, we construct generalized entropy weights that simultaneously enforce a balancing constraint based on outcome-related information and a debiasing constraint tied to a working propensity-score model. This yields a common inferential template for parameters defined by general estimating equations under missing at random, spanning causal inference, semi-supervised learning, and regression with missing covariates.
The framework has several important consequences. When either the outcome-regression model or the propensity-score model is correctly specified, the resulting estimator is consistent, and when both are correct it attains the semiparametric efficiency bound. More distinctively, Corollary 3 shows that under correct propensity-score specification the proposed estimator is never less efficient than classical AIPW and is strictly more efficient whenever the debiasing information is not already contained in the augmentation space. In this sense, generalized entropy calibration does more than stabilize weights: it enlarges the effective projection space used for augmentation and thereby improves efficiency under outcome-model misspecification.
The simulations and real-data analyses support this theoretical picture. Across the examples considered in the paper, the proposed estimators recover the efficient benchmark when both working models are correct, retain the expected robustness when only one working model is correct, and show their clearest gains when the propensity-score model is correct but the outcome model is misspecified. The empirical results therefore reinforce the main theoretical message of the paper: calibration can serve not only as a balancing device, but also as a general mechanism for doubly robust and efficient inference with partially observed data.
Several extensions merit further investigation. First, the current framework assumes a missing-at-random mechanism; extending the calibration formulation to nonignorable missingness would substantially broaden its scope. Second, when the calibration function is high-dimensional relative to the sample size, the entropy optimization problem may require regularization or dimension reduction for stable implementation. More broadly, the proposed framework suggests that calibration can provide a common language for weighting, augmentation, and prediction in a wide range of problems involving incomplete or integrated data.
{acks}
[Data and Code Availability] The real data analyses in this study use publicly available data. The National Supported Work (NSW) experimental data (LaLonde, 1986; Dehejia and Wahba, 1999) are available from the National Bureau of Economic Research (NBER) at http://www.nber.org/~rdehejia/data/ and can be loaded using the haven R package. The PSID and CPS comparison samples are obtained from the DAAG R package. The NHANES data are publicly available from the Centers for Disease Control and Prevention (CDC) at https://www.cdc.gov/nchs/nhanes and can be accessed using the nhanesA R package. All simulation and analysis code used to generate the results in this article is available from the authors upon request.
Disclosure Statement
No potential conflict of interest was reported by the author(s).